I have to create one statement in english representing an algebraic relationship?
How? I don't understand what am i suppost to do and in the book it doesn't explain it enough.can anyone give me examples or help me through it.

Answer by venugopalramana(3286)

I have to create one statement in english representing an algebraic relationship?
How? I don't understand what am i suppost to do and in the book it doesn't explain it enough.can anyone give me examples or help me through it.
LET US SAY JOHN IS ELDER TO PETER BY 3 YEARS.
SO IF PETER IS 10 YRS,JOHN IS 10+3= 13 YRS....
IF PETER IS 5 YRS,JOHN IS 5+3=8 YRS....ETC...
AS WE ARE NOT GIVEN THE ACTUAL AGES,WE TAKE
LET PETER BE P YEARS (IF YOU HAVE ANY INHIBITION OR DOUBT WITH USE OF A SYMBOL , JUST IMAGINE IT TO BE EQUAL TO SOME 10 AS ABOVE AND PROCEED.FINALLY PUT P IN PLACE OF 10)....SO JOHN IS P+3 YEARS....IF WE CALL JOHN'S AGE AS J YEARS THEN WE GET AN EQUATION
J=P+3
THIS IS A SIMPLE EXAMPLE.PROCEED ON SIMILAR LINES TO GET BIGGER EXPRESSIONS.WHEN IN TROUBLE DO IT IN SIMPLE NUMBERS AND REPLACE THOSE NUMBERS BY SYMBOLS FINALLY.
THIS TYPE OF EXPRESSION IN SYMBOLS IS CALLED ALGEBRAIC RELATION.